### v1

# A Closed-form Solution to the Wahba Problem for Pairwise Similar Quaternions

Abstract.
We present a closed-form solution to Wahba’s problem in the quaternion domain for the special case of two vector observations. Existing approaches, including Davenport’s q{q}-method, QUEST, Horn’s method, and ESOQ algorithms, recover the optimal quaternion through the eigendecomposition of a $4\times 4$ matrix or iterative numerical methods. Consequently, these methods do not reveal the analytic structure of the optimal quaternion.
In this work, we derive an explicit analytical characterization of all quaternions that yield zero Wahba cost for the case $\ell=2$. Our approach builds on a connection between quaternion similarity, the singular Sylvester equation aq=qb{a}{q}={q}{b}, and quaternion square roots established in our previous work [1]. We provide (i) necessary and sufficient conditions under which the Wahba’s cost function is zero and (ii) a closed-form parameterization of all such quaternions. This eliminates the need for eigenvalue computations and enables a direct algebraic understanding of the underlying geometry of Wahba’s problem.
1. Introduction
Determining the relative orientation between two Cartesian coordinate systems is a fundamental in many practical applications, such as spacecraft attitude estimation, robotics, and autonomous driving. These applications require fast and accurate information about the object’s orientation. Solutions often rely on vector observations expressed in two coordinate systems. The problem of finding the relationship between two sets of vector measurements is formally known as the Wahba’s problem [2]. Wahba’s problem has been addressed using various mathematical formulations, including rotation matrices, direction cosine matrices, and quaternions. Among these, quaternion-based approaches are particularly effective due to their compact representation, absence of singularities, and computational efficiency.
Wahba’s problem in quaternions states: Given two sets of pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{\ell}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{\ell}\}$, where $\ell\geq 2$, Wahba’s problem aims at finding a nonzero quaternion q{q} that minimizes the cost function
(1.1) $f({q})=\sum_{\ell=1}^{n}\mathopen{\lvert}{q}^{-1}{a}_{\ell}{q}-{b}_{\ell}\mathclose{\rvert}^{2}.$
Pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{\ell}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{\ell}\}$ represent vectors observed in the sensor frames $\mathcal{\!}{A}$ and $\mathcal{\!}{B}$, respectively.
Several well-established algorithms solve Wahba’s problem, including Davenport’s q{q}-method [3], the QUEST algorithms [4, 5, 6], Horn’s formulation [7], and ESOQ methods [8, 9] for the general case $\ell\geq 2$. All of these methods obtain the optimal quaternion either through the eigendecomposition of a $4\times 4$ matrix or through iterative numerical procedures. As a result, the solution is obtained indirectly via numerical linear algebra, without exposing the analytic relationship between the structure of Wahba’s cost function and the form of its optimal quaternion. Consequently, to the best of our knowledge, the literature contains no closed-form quaternion solution that fully exploits the special structure of the case $\ell=2$.
In this paper, we solve the Wahba’s problem for the special and practically important case $\ell=2$. We establish a direct relation between the quaternionic linear equation aqqb=0{a}{q}-{q}{b}=0, commonly known as the Sylvester equation, studied in our previous work [1], and the Wahba’s cost function in (1.1). This relation leads to the two main contributions of this work: (i) the necessary and sufficient condition for the existence of solutions and (ii) an analytical expression for the solution set of f(q)=0f({q})=0 when $\ell=2$. Our approach avoids conventional matrix-based techniques, such as iterative eigenvalue computation or matrix inversion, thereby improving computational efficiency and analytical clarity of the solution.
The remainder of the paper is structured as follows. Section 2 reviews the necessary quaternionic preliminaries, introduces a notion of pairwise similarity, and summarizes background results from [1]. Section 4 presents the main results: the existence criterion and the complete closed-form solution set of f(q)=0f({q})=0 when $\ell=2$. Section 5 concludes the paper.
2. Preliminaries
This section introduces the quaternion notation and reviews the relevant algebraic properties of quaternions used throughout the paper. In particular, the notion of pairwise quaternion similarity is introduced, which will be the central concept governing the solvability of Wahba’s cost function for $\ell=2$.
[…]
${a}{b}\coloneqq(\mathrm{Re}\,{a}{})(\mathrm{Re}\,{b}{})-{\mathrm{Im}\,{a}}\!\cdot{\mathrm{Im}\,{b}}+(\mathrm{Re}\,{a}{})({\mathrm{Im}\,{b}})+(\mathrm{Re}\,{b}{})({\mathrm{Im}\,{a}})+{\mathrm{Im}\,{a}}\times{\mathrm{Im}\,{b}}.$
where (,×)($\cdot\,$$\times$) denote the scalar and cross product, respectively. It also holds that
(2.1) $\mathopen{\lvert}{a}{b}\mathclose{\rvert}=\mathopen{\lvert}{a}\mathclose{\rvert}\kern 1.00006pt\mathopen{\lvert}{b}\mathclose{\rvert},\qquad\mathrm{Re}\,({a}{b})=\mathrm{Re}\,({b}{a}).$
For pure quaternions a,b{a},{b}, the following relations hold
${a}^{2}=-{a}\cdot{a}=-\mathopen{\lvert}{a}\mathclose{\rvert}^{2},\quad{a}{b}=-{a}\cdot{b}+{a}\times{b},\quad{a}{b}+{b}{a}=2\mathrm{Re}\,({a}{b}).$
Norm of the sum of two quaternions a{a}, b{b} is given as [10]
(2.2) $\mathopen{\lvert}{a}\pm{b}\mathclose{\rvert}^{2}=\mathopen{\lvert}{a}\mathclose{\rvert}^{2}\pm 2\mathrm{Re}\,({a}{b}^{*})+\mathopen{\lvert}{b}\mathclose{\rvert}^{2}.$
Lemma 2.1.
Let a,b{a},{b} be nonreal pure quaternions and define ${c}=\mathrm{Im}\,({a}{b})={a}\times{b}$. Then, ac=ca{a}{c}=-{c}{a}.
Proof.
Since ${c}={a}\times{b}$, we have ${a}\cdot{c}=0$, and hence ac=ca{a}{c}=-{c}{a}. ∎
Definition 2.2.
Nonreal quaternions a,b{a},{b} are called similar if there exists a nonzero quaternion p{p} such that ${p}^{-1}{a}{p}={b}$. In that case we write ${a}\sim{b}$.
Remark 2.3.
In Definition 2.2, similarity was introduced only for nonreal quaternions. More generally, the notion can also be extended to real quaternions. However, throughout this work, we consider only the nonreal quaternions ${a}_{\ell},{b}_{\ell}$. Otherwise, ${p}^{-1}{a}{p}-{b}={a}-{b}$, which is independent of q{q}.
Lemma 2.4 ([1]).
Nonreal quaternions a,b{a},{b} are similar if and only if $\mathrm{Re}\,{a}{}=\mathrm{Re}\,{b}{}$ and $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$.
Definition 2.5.
Two pairs of nonreal quaternions $({a}_{1},{a}_{2})$ and $({b}_{1},{b}_{2})$ are called pairwise similar if there exists a nonzero quaternion p{p} such that ${p}^{-1}{a}_{1}{p}={b}_{1}$ and ${p}^{-1}{a}_{2}{p}={b}_{2}$. In that case we write $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$.
Lemma 2.6.
For nonreal quaternions ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ it holds $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$ if and only if ${a}_{1}\sim{b}_{1}$ and ${a}_{2}\sim{b}_{2}$ and $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$.
Proof.
Assume first that $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$. Then, by Definition 2.5, there exists a nonzero quaternion p{p} such that ${p}^{-1}{a}_{1}{p}={b}_{1}$ and ${p}^{-1}{a}_{2}{p}={b}_{2}$. It follows that ${a}_{1}\sim{b}_{1}$ and ${a}_{2}\sim{b}_{2}$. Furthermore, adding the two equalities yields
${p}^{-1}({a}_{1}+{a}_{2}){p}={b}_{1}+{b}_{2}\quad\Longrightarrow\quad({a}_{1}+{a}_{2})\sim({b}_{1}+{b}_{2}).$
Then, by Lemma 2.4 and the property in (2.2) we obtain
(2.3) $\mathopen{\lvert}{a}_{1}\mathclose{\rvert}^{2}+2\mathrm{Re}\,({a}_{1}{a}_{2}^{*})+\mathopen{\lvert}{a}_{2}\mathclose{\rvert}^{2}=\mathopen{\lvert}{a}_{1}+{a}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{1}+{b}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{1}\mathclose{\rvert}^{2}+2\mathrm{Re}\,({b}_{1}{b}_{2}^{*})+\mathopen{\lvert}{b}_{2}\mathclose{\rvert}^{2}.$
With $\mathopen{\lvert}{a}_{1}\mathclose{\rvert}=\mathopen{\lvert}{b}_{1}\mathclose{\rvert}$ and $\mathopen{\lvert}{a}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{2}\mathclose{\rvert}$, it equivalently holds $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$. Conversely, assume ${a}_{1}\sim{b}_{1}$ and ${a}_{2}\sim{b}_{2}$ and $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$. Then it holds $\mathrm{Re}\,({a}_{1}+{a}_{2})=\mathrm{Re}\,({b}_{1}+{b}_{2})$. Furthermore, by property in (2.3) we have $\mathopen{\lvert}{a}_{1}+{a}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{1}+{b}_{2}\mathclose{\rvert}$. Hence, $({a}_{1}+{a}_{2})\sim({b}_{1}+{b}_{2})$, that is there exists a nonzero quaternion q{q} such that
(2.4) ${q}^{-1}({a}_{1}+{a}_{2}){q}={b}_{1}+{b}_{2}.$
Assume that ${q}_{1}$ is a nonzero quaternion with ${q}_{1}^{-1}{a}_{1}{q}_{1}={b}_{1}$. Set ${q}={q}_{1}{q}_{2}$, where ${q}_{2}$ is a nonzero quaternion such that ${b}_{1}{q}_{2}={q}_{2}{b}_{1}$. Then we have
(2.5) ${q}^{-1}{a}_{1}{q}={q}_{2}^{-1}{q}_{1}^{-1}{a}_{1}{q}_{1}{q}_{2}={q}_{2}^{-1}{b}_{1}{q}_{2}={q}_{2}^{-1}{q}_{2}{b}_{1}={b}_{1}.$
Substituting (2.5) into (2.4) yields ${q}^{-1}{a}_{2}{q}={b}_{2}$, which completes the proof. ∎
3. Singular Sylvester equation in quaternions
In this section, we summarize the results from [1] concerning the homogeneous singular Sylvester equation in quaternions
(3.1) ${a}{q}-{q}{b}=0$
and its relation to quaternion square roots. We recall the necessary and sufficient condition for solvability and the general solution of (3.1), which sets the foundation for the results developed in Section 4.
Lemma 3.1 ([1]).
Let a{a} be a nonzero quaternion. Then, ${q}^{2}-{a}=0$ has the solutions
${q}=\pm\sqrt{\mathopen{\lvert}{a}\mathclose{\rvert}}\,\frac{{p}}{\mathopen{\lvert}{p}\mathclose{\rvert}}$,
where p{p} is either an arbitrary pure quaternion if a{a} is a negative real number or ${p}={a}+\mathopen{\lvert}{a}\mathclose{\rvert}$ if a{a} is a nonreal quaternion or a positive real number.
Lemma 3.2 ([1]).
Let a,b{a},{b} be nonreal quaternions. Then, (3.1) has a nonzero solution if and only if ${a}\sim{b}$.
Theorem 3.3 ([1]).
Let a,b{a},{b} be similar quaternions and $\lambda,\mu$ arbitrary real numbers with $\mathopen{\lvert}\lambda\mathclose{\rvert}+\mathopen{\lvert}\mu\,\mathrm{Im}\,{({a}+{b})}\mathclose{\rvert}\neq 0$. Then, the nonzero solutions to (3.1) are given as
(3.2) ${q}=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\,\mathrm{Im}\,{({a}+{b})},$
subject to ${\mathrm{Im}\,{a}}\cdot{\mathrm{Im}\,{q}}=0$ if ${\mathrm{Im}\,{a}}=-{\mathrm{Im}\,{b}}$.
4. Solutions to Wahba’s problem in quaternions
We consider Wahba’s problem for the case of two vector observations such that
(4.1) $f({q})=\sum_{\ell=1}^{2}\mathopen{\lvert}{q}^{-1}{a}_{\ell}{q}-{b}_{\ell}\mathclose{\rvert}^{2}=0.$
In this section, we establish the connection between the solvability of (4.1) and quaternion pairwise similarity. We show that (4.1) corresponds to pairs of singular Sylvester equations, which underpins the algebraic structure of the Wahba’s problem for the special case $\ell=2$. Based on this, we show the necessary and sufficient conditions for existance of solution to (4.1) and derive their explicit analytical representation.
Lemma 4.1.
Let a,b{a},{b} be nonreal quaternions. Then, $\mathopen{\lvert}{q}^{-1}{a}{q}-{b}\mathclose{\rvert}^{2}=0$ has a solution if and only if ${a}\sim{b}$.
Proof.
Let p{p} be a nonzero quaternion such that
${p}^{-1}{a}{p}={b}\quad\Longleftrightarrow\quad{p}^{-1}{a}{p}-{b}=0\quad\Longleftrightarrow\quad\mathopen{\lvert}{p}^{-1}{a}{p}-{b}\mathclose{\rvert}^{2}=0.$
Hence, q=p{q}={p} satisfies the claim. ∎
Corollary 4.1.1.
Let a,b{a},{b} be nonreal quaternions with ${a}\sim{b}$. Then, the general solution of $\mathopen{\lvert}{q}^{-1}{a}{q}-{b}\mathclose{\rvert}^{2}=0$ is given in (3.2).
Proof.
Let q{q} be a nonzero quaternion such that $\mathopen{\lvert}{q}^{-1}{a}{q}-{b}\mathclose{\rvert}^{2}=0$. Then, it equivalently holds ${q}^{-1}{a}{q}-{b}=0$ and equivalently aqqb=0{a}{q}-{q}{b}=0. By applying Theorem 3.3, we obtain (3.2). Hence, the set of all nonzero solutions of the singular Sylvester equation aqqb=0{a}{q}-{q}{b}=0 coincides with the set of all solutions of $\mathopen{\lvert}{q}^{-1}{a}{q}-{b}\mathclose{\rvert}^{2}=0$. ∎
Corollary 4.1.2.
Let a,b{a},{b} be nonreal quaternions with ${a}\sim{b}$. Then, for all nonzero quaternions q{q} it holds
(4.2) $f({q})=\mathopen{\lvert}{q}^{-1}{a}{q}-{b}\mathclose{\rvert}^{2}=\mathopen{\lvert}{q}^{-1}({\mathrm{Im}\,{a}}){q}-{\mathrm{Im}\,{b}}\mathclose{\rvert}^{2}.$
Proof.
Since $\mathrm{Re}\,{b}{}=\mathrm{Re}\,{a}{}=(\mathrm{Re}\,{a}{}){q}^{-1}{q}={q}^{-1}(\mathrm{Re}\,{a}{}){q}$, we have
$\qquad{q}^{-1}{a}{q}-{b}={q}^{-1}(\mathrm{Re}\,{a}{}+{\mathrm{Im}\,{a}}){q}-(\mathrm{Re}\,{b}{}+{\mathrm{Im}\,{b}})={q}^{-1}({\mathrm{Im}\,{a}}){q}-{\mathrm{Im}\,{b}}.\qquad\qquad\qed$
When equation (4.1) is considered, the quaternions ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ may, by property (4.2) and without loss of generality, be considered as pure. We will make use of this assumption in proofs of the following theorems. However, the statements of the theorems will be presented in their general form.
Theorem 4.2.
Let ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ be nonreal quaternions. Then, (4.1) has a solution if and only if $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$.
Proof.
Assume that (4.1) holds. Then, there exists a nonzero quaternion q{q} such that ${q}^{-1}{a}_{1}{q}={b}_{1}$ and ${q}^{-1}{a}_{2}{q}={b}_{2}$. By Definition 2.5 it holds $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$. Conversely, assume $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$. Then, there exists a nonzero quaternion p{p} such that ${p}^{-1}{a}_{1}{p}={b}_{1}$ and ${p}^{-1}{a}_{2}{p}={b}_{2}$. Hence, q=p{q}={p} is a solution of (4.1). ∎
Lemma 4.3.
Let ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ be nonreal quaternions with $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$. Then, (4.1) is equivalent to
(4.3) $\mathopen{\lvert}{q}^{-1}({\mathrm{Im}\,{a}}_{1}){q}-{\mathrm{Im}\,{b}}_{1}\mathclose{\rvert}^{2}+\mathopen{\lvert}{q}^{-1}({\mathrm{Im}\,{a}}_{1}\times{\mathrm{Im}\,{a}}_{2}){q}-{\mathrm{Im}\,{b}}_{1}\times{\mathrm{Im}\,{b}}_{2}\mathclose{\rvert}^{2}=0.$
Proof.
By Corollary 4.1.2, we may, without loss of generality, consider ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ as pure. Let q{q} be a solution of (4.1). Then, it holds
$f({q})=0\quad\Longleftrightarrow\quad{q}^{-1}{a}_{1}{q}={b}_{1}\quad\text{and}\quad{q}^{-1}{a}_{2}{q}={b}_{2}.$
Multiplying ${q}^{-1}{a}_{2}{q}={b}_{2}$ by ${q}^{-1}{a}_{1}{q}$ from the left and with $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$, we obtain
${q}^{-1}{a}_{1}{a}_{2}{q}={q}^{-1}{a}_{1}{q}{b}_{2}={b}_{1}{b}_{2}\quad\Longleftrightarrow\quad{q}^{-1}[\mathrm{Im}\,({a}_{1}{a}_{2})]{q}=\mathrm{Im}\,({b}_{1}{b}_{2}),$
and hence (4.3), which completes the proof. ∎
Theorem 4.4.
Let ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ be nonreal quaternions with $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$ and $\lambda_{1},\lambda_{2}$, and $\mu$ real numbers with $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu_{1}\,\mathrm{Im}\,{({a}_{1}+{b}_{1})}\mathclose{\rvert}\neq 0$ and $\lambda_{2}\neq 0$. Then, the general solution to (4.1) is given as ${q}={q}_{1}{q}_{2}$, where
(4.4) ${q}_{1}$ $=\lambda_{1}\sqrt{({\mathrm{Im}\,{a}}_{1})({\mathrm{Im}\,{b}}_{1}^{*})}+\mu_{1}\,\mathrm{Im}\,{({a}_{1}+{b}_{1})},$
(4.5) ${q}_{2}$ $=\lambda_{2}\sqrt{{q}_{1}^{*}({\mathrm{Im}\,{a}}_{1}\times{\mathrm{Im}\,{a}}_{2}){q}_{1}({\mathrm{Im}\,{b}}_{2}\times{\mathrm{Im}\,{b}}_{1})},$
subject to
${\mathrm{Im}\,{a}}_{1}\cdot{\mathrm{Im}\,{q}}_{1}=0\quad$ $\text{if}\quad{\mathrm{Im}\,{a}}_{1}=-{\mathrm{Im}\,{b}}_{1},$
${q}_{1}^{*}({\mathrm{Im}\,{a}}_{1}\times{\mathrm{Im}\,{a}}_{2}){q}_{1}\cdot{\mathrm{Im}\,{q}}_{2}=0\quad$ $\text{if}\quad{q}_{1}^{*}({\mathrm{Im}\,{a}}_{1}\times{\mathrm{Im}\,{a}}_{2}){q}_{1}=-{\mathrm{Im}\,{b}}_{1}\times{\mathrm{Im}\,{b}}_{2}.$
Proof.
By Corollary 4.1.2, we may, without loss of generality, consider ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ as pure. Let q{q} be a solution of (4.1). Then, by Lemma 4.3, it equivalently holds
$\mathopen{\lvert}{q}^{-1}{a}_{1}{q}-{b}_{1}\mathclose{\rvert}^{2}+\mathopen{\lvert}{q}^{-1}({a}_{1}\times{a}_{2}){q}-{b}_{1}\times{b}_{2}\mathclose{\rvert}^{2}=0$
and equivalently
(4.6) $\mathopen{\lvert}{q}^{-1}{a}_{1}{q}-{b}_{1}\mathclose{\rvert}^{2}=0\quad\text{and}\quad\mathopen{\lvert}{q}^{-1}({a}_{1}\times{a}_{2}){q}-{b}_{1}\times{b}_{2}\mathclose{\rvert}^{2}=0.$
Let ${q}_{1}$ be a nonzero quaternion such that $\mathopen{\lvert}{q}_{1}^{-1}{a}_{1}{q}_{1}-{b}_{1}\mathclose{\rvert}^{2}=0$. Then, it equivalently holds ${a}_{1}{q}_{1}={q}_{1}{b}_{1}$. Moreover, by Corollary 4.1.1 we obtain ${q}_{1}$ in (4.4) subject to $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu\,({a}_{1}+{b}_{1})\mathclose{\rvert}\neq 0$ and ${a}_{1}\cdot{\mathrm{Im}\,{q}}_{1}=0$ if ${a}_{1}=-{b}_{1}$. For any nonzero quaternion q{q} there exists a unique nonzero quaternion ${q}_{2}$ such that ${q}={q}_{1}{q}_{2}$. Then, substituting ${q}={q}_{1}{q}_{2}$ and ${a}_{1}{q}_{1}={q}_{1}{b}_{1}$ into the left side equality in (4.6) yields
(4.7) ${b}_{1}=({q}_{1}{q}_{2})^{-1}{a}_{1}{q}_{1}{q}_{2}={q}_{2}^{-1}{q}_{1}^{-1}{q}_{1}{b}_{1}{q}_{2}={q}_{2}^{-1}{b}_{1}{q}_{2}\quad\Longleftrightarrow\quad{q}_{2}{b}_{1}={b}_{1}{q}_{2}.$
Furthermore, substituting ${q}={q}_{1}{q}_{2}$ into the right side equality in (4.6) yields
(4.8) $({q}_{1}{q}_{2})^{-1}({a}_{1}\times{a}_{2}){q}_{1}{q}_{2}-{b}_{1}\times{b}_{2}={q}_{2}^{-1}{q}_{1}^{-1}({a}_{1}\times{a}_{2}){q}_{1}{q}_{2}-{b}_{1}\times{b}_{2}=0.$
Define ${a}_{3}={q}_{1}^{-1}({a}_{1}\times{a}_{2}){q}_{1}$ and ${b}_{3}={b}_{1}\times{b}_{2}$. Then, by the left side property in (2.1), and since $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$, we obtain
$\mathopen{\lvert}{a}_{3}\mathclose{\rvert}=\mathopen{\lvert}{q}_{1}^{-1}({a}_{1}\times{a}_{2}){q}_{1}\mathclose{\rvert}=\mathopen{\lvert}{a}_{1}\times{a}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{1}\times{b}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{3}\mathclose{\rvert}$,
and with $\mathrm{Re}\,{a}_{3}=\mathrm{Re}\,{b}_{3}=0$ further ${a}_{3}\sim{b}_{3}$. Moreover, by Lemma 2.1 it holds $({a}_{1}\times{a}_{2}){a}_{1}=-{a}_{1}({a}_{1}\times{a}_{2})$ and ${b}_{3}{b}_{1}=-{b}_{1}{b}_{3}$. Furthermore, by the left side equality in (4.6) we have ${b}_{1}={q}^{-1}{a}_{1}{q}$ and hence
(4.9) ${a}_{3}{b}_{1}={q}^{-1}({a}_{1}\times{a}_{2}){q}{q}^{-1}{a}_{1}{q}=-{q}^{-1}{a}_{1}{q}{q}^{-1}({a}_{1}\times{a}_{2}){q}=-{b}_{1}{a}_{3}.$
Now, multiplying (4.8) by ${q}_{2}$ from the left yields ${a}_{3}{q}_{2}-{q}_{2}{b}_{3}=0$ such that by Theorem 3.3 and for real numbers $\lambda_{2},\mu_{2}$ with $\mathopen{\lvert}\lambda_{2}\mathclose{\rvert}+\mathopen{\lvert}\mu_{2}\,({a}_{3}+{b}_{3})\mathclose{\rvert}\neq 0$ we obtain
(4.10) ${q}_{2}=\lambda_{2}\sqrt{{a}_{3}{b}_{3}^{*}}+\mu_{2}({a}_{3}+{b}_{3}),$
subject to ${a}_{3}\cdot{\mathrm{Im}\,{q}}_{2}=0$ if ${a}_{3}=-{b}_{3}$. Substituting  (4.10) into the right side equality in (4.7) and with (4.9) and ${b}_{3}{b}_{1}=-{b}_{1}{b}_{3}$ we obtain
$0={q}_{2}{b}_{1}-{b}_{1}{q}_{2}=\mu_{2}({a}_{3}{b}_{1}+{b}_{3}{b}_{1}-{b}_{1}{a}_{3}-{b}_{1}{b}_{3})=2\mu_{2}({a}_{3}+{b}_{3}){b}_{1}$
and hence $\mu_{2}({a}_{3}+{b}_{3})=0$, which completes the proof. ∎
5. Conclusion
This work establishes the necessary and sufficient condition under which Wahba’s cost function for two vector observations equals zero. Subsequently, we derived the corresponding closed-form solution for pairwise similar quaternions, by revealing the connection between the singular Sylvester equation in quaternions and quaternionic square roots. We obtained a direct analytical pathway to the closed-form solution in the two-observation case. The presented results offer a complete theoretical characterisation of the Wahba problem in quaternions and lay a rigorous foundation for further generalisations in quaternion-based orientation estimation.
Acknowledgement
This work was supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden (Project ID: 390696704) and by the Federal Ministry of Education and Research of Germany in the programme “Souverän. Digital. Vernetzt.” Joint project 6G-life (Project ID: 16KISK001K).

### v2

@@
-A Closed-form Solution to the Wahba Problem for Pairwise Similar Quaternions
+A Closed-form Solution to the Wahba’s Problem for Pairwise Similar Quaternions
 Abstract.
@@
 Consequently, these methods do not reveal the analytic structure of the optimal quaternion.
-In this work, we derive an explicit analytical characterization of all quaternions that yield zero Wahba cost for the case $\ell=2$.
+In this work, we derive an explicit analytical characterization of all quaternions that yield zero Wahba’s cost for the case $n=2$.
 Our approach builds on a connection between quaternion similarity, the singular Sylvester equation aq=qb{a}{q}={q}{b}, and quaternion square roots established in our previous work [5].
@@
 This eliminates the need for eigenvalue computations and enables a direct algebraic understanding of the underlying geometry of Wahba’s problem.
+Keywords: Quaternions; Wahba’s problem; Similarity; Pairwise similarity; Closed-form solution
+Mathematical sub classification numbers (MSC) 2020: 15A24, 15B33
 1. Introduction
@@
 Solutions often rely on vector observations expressed in two coordinate systems.
-The problem of finding the relationship between two sets of vector measurements is formally known as the Wahba’s problem [2].
+The problem of finding the relationship between two sets of vector measurements is formally known as Wahba’s problem [10].
 Wahba’s problem has been addressed using various mathematical formulations, including rotation matrices, direction cosine matrices, and quaternions.
 Among these, quaternion-based approaches are particularly effective due to their compact representation, absence of singularities, and computational efficiency.
-Wahba’s problem in quaternions states: Given two sets of pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{\ell}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{\ell}\}$, where $\ell\geq 2$, Wahba’s problem aims at finding a nonzero quaternion q{q} that minimizes the cost function
+Wahba’s problem in quaternions states: Given two sets of pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{n}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{n}\}$, where $n\geq 2$, Wahba’s problem aims at finding a nonzero quaternion q{q} that minimizes the cost function
 (1.1) $f({q})=\sum_{\ell=1}^{n}\mathopen{\lvert}{q}^{-1}{a}_{\ell}{q}-{b}_{\ell}\mathclose{\rvert}^{2}.$
-Pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{\ell}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{\ell}\}$ represent vectors observed in the sensor frames $\mathcal{\!}{A}$ and $\mathcal{\!}{B}$, respectively.
-Several well-established algorithms solve Wahba’s problem, including Davenport’s q{q}-method [3], the QUEST algorithms [4, 5, 6], Horn’s formulation [7], and ESOQ methods [8, 9] for the general case $\ell\geq 2$.
+Pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{n}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{n}\}$ represent vectors observed in the sensor frames $\mathcal{\!}{A}$ and $\mathcal{\!}{B}$, respectively.
+Several well-established algorithms solve Wahba’s problem, including Davenport’s q{q}-method [1], the QUEST algorithms [8, 9, 7], Horn’s formulation [2], and ESOQ methods [4, 3] for the general case $n\geq 2$.
 All of these methods obtain the optimal quaternion either through the eigendecomposition of a $4\times 4$ matrix or through iterative numerical procedures.
 As a result, the solution is obtained indirectly via numerical linear algebra, without exposing the analytic relationship between the structure of Wahba’s cost function and the form of its optimal quaternion.
-Consequently, to the best of our knowledge, the literature contains no closed-form quaternion solution that fully exploits the special structure of the case $\ell=2$.
-In this paper, we solve the Wahba’s problem for the special and practically important case $\ell=2$.
+Consequently, to the best of our knowledge, the literature contains no closed-form quaternion solution that fully exploits the special structure of the case $n=2$.
+In this paper, we solve the Wahba’s problem for the special and practically important case $n=2$.
 We establish a direct relation between the quaternionic linear equation aqqb=0{a}{q}-{q}{b}=0, commonly known as the Sylvester equation, studied in our previous work [5], and the Wahba’s cost function in (1.1).
-This relation leads to the two main contributions of this work: (i) the necessary and sufficient condition for the existence of solutions and (ii) an analytical expression for the solution set of f(q)=0f({q})=0 when $\ell=2$.
+This relation leads to the two main contributions of this work: (i) the necessary and sufficient condition for the existence of solutions and (ii) an analytical expression for the solution set of f(q)=0f({q})=0 when $n=2$.
 Our approach avoids conventional matrix-based techniques, such as iterative eigenvalue computation or matrix inversion, thereby improving computational efficiency and analytical clarity of the solution.
@@
 Section 2 reviews the necessary quaternionic preliminaries, introduces a notion of pairwise similarity, and summarizes background results from [5].
-Section 4 presents the main results: the existence criterion and the complete closed-form solution set of f(q)=0f({q})=0 when $\ell=2$.
+Section 4 presents the main results: the existence criterion and the complete closed-form solution set of f(q)=0f({q})=0 when $n=2$.
 Section 5 concludes the paper.
@@
 This section introduces the quaternion notation and reviews the relevant algebraic properties of quaternions used throughout the paper.
-In particular, the notion of pairwise quaternion similarity is introduced, which will be the central concept governing the solvability of Wahba’s cost function for $\ell=2$.
+In particular, the notion of pairwise quaternion similarity is introduced, which will be the central concept governing the solvability of Wahba’s cost function for $n=2$.
 […]
@@
 In this section, we establish the connection between the solvability of (4.1) and quaternion pairwise similarity.
-We show that (4.1) corresponds to pairs of singular Sylvester equations, which underpins the algebraic structure of the Wahba’s problem for the special case $\ell=2$.
+We show that (4.1) corresponds to pairs of singular Sylvester equations, which underpins the algebraic structure of the Wahba’s problem for the special case $n=2$.
 Based on this, we show the necessary and sufficient conditions for existance of solution to (4.1) and derive their explicit analytical representation.
@@
 Theorem 4.4.
-Let ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ be nonreal quaternions with $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$ and $\lambda_{1},\lambda_{2}$, and $\mu$ real numbers with $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu_{1}\,\mathrm{Im}\,{({a}_{1}+{b}_{1})}\mathclose{\rvert}\neq 0$ and $\lambda_{2}\neq 0$.
+Let ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ be nonreal quaternions with $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$ and $\lambda_{1},\lambda_{2}$, and $\mu_{1}$ real numbers with $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu_{1}\,\mathrm{Im}\,{({a}_{1}+{b}_{1})}\mathclose{\rvert}\neq 0$ and $\lambda_{2}\neq 0$.
 Then, the general solution to (4.1) is given as ${q}={q}_{1}{q}_{2}$, where
@@
 Then, it equivalently holds ${a}_{1}{q}_{1}={q}_{1}{b}_{1}$.
-Moreover, by Corollary 4.1.1 we obtain ${q}_{1}$ in (4.4) subject to $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu\,({a}_{1}+{b}_{1})\mathclose{\rvert}\neq 0$ and ${a}_{1}\cdot{\mathrm{Im}\,{q}}_{1}=0$ if ${a}_{1}=-{b}_{1}$.
+Moreover, by Corollary 4.1.1 we obtain ${q}_{1}$ in (4.4) subject to $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu_{1}({a}_{1}+{b}_{1})\mathclose{\rvert}\neq 0$ and ${a}_{1}\cdot{\mathrm{Im}\,{q}}_{1}=0$ if ${a}_{1}=-{b}_{1}$.
 For any nonzero quaternion q{q} there exists a unique nonzero quaternion ${q}_{2}$ such that ${q}={q}_{1}{q}_{2}$.
@@
 Define ${a}_{3}={q}_{1}^{-1}({a}_{1}\times{a}_{2}){q}_{1}$ and ${b}_{3}={b}_{1}\times{b}_{2}$.
-Then, by the left side property in (2.1), and since $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$, we obtain
+Then, by the left side property in (2.1) and since $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$, we obtain
 $\mathopen{\lvert}{a}_{3}\mathclose{\rvert}=\mathopen{\lvert}{q}_{1}^{-1}({a}_{1}\times{a}_{2}){q}_{1}\mathclose{\rvert}=\mathopen{\lvert}{a}_{1}\times{a}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{1}\times{b}_{2}\mathclose{\rvert}=\mathopen{\lvert}{b}_{3}\mathclose{\rvert}$,
@@
 This work establishes the necessary and sufficient condition under which Wahba’s cost function for two vector observations equals zero.
-Subsequently, we derived the corresponding closed-form solution for pairwise similar quaternions, by revealing the connection between the singular Sylvester equation in quaternions and quaternionic square roots.
+Subsequently, we derived the corresponding closed-form solution for pairwise similar quaternions by revealing the connection between the singular Sylvester equation in quaternions and quaternionic square roots.
 We obtained a direct analytical pathway to the closed-form solution in the two-observation case.
-The presented results offer a complete theoretical characterisation of the Wahba problem in quaternions and lay a rigorous foundation for further generalisations in quaternion-based orientation estimation.
+The presented results offer a complete theoretical characterisation of the Wahba’s problem in quaternions and lay a rigorous foundation for further generalisations in quaternion-based orientation estimation.
 Acknowledgement